The measure of the first angle of a triangle is four times as large as the measure of the second angle. The third angle measures 15 degrees less than twice the first angle. Find the measures of the angles.
Answer by rfer(16322)

4x+x+2(4x)-15=180
4x+x+8x-15=180
13x=195
x=195/13
x=15
4x=60
2(4x)-15=105